# A Survey on Self-Supervised Learning for Non-Sequential Tabular Data

### v1

@@ 6.2 Cross-Table Transferability
Directly learning representations from a table requires a trained model per downstream dataset and suffers from strict features between training and testing data, costing a severe burden in scaling and transferring to various problems. Therefore, how to learn representations across tables has been a critical demand for NS-TD systems in real-world scenarios, which benefits from reducing efforts on engineering features based on each dataset. Recent approaches achieving transferability involve SSL pre-training from PLMs to contextualize knowledge with coherent semantics (e.g., LIFT, TP-BERTa, GReaT) and from scratch with fine-grained feature encodings (e.g., TransTab, XTab, UniTabE). These methods have demonstrated that pre-training with SSL4NS-TD confers advantages for adaptation to incremental columns, low-resource scenarios, and missing value predictions Yang et al. (2023); Zhu et al. (2023).
@@ 6.3 Domain Knowledge Integration
Tabular applications often have the need to incorporate expert knowledge to infer the results (e.g., clinical trials in the medical and real estate market in finance). Du et al. (2023) discovered that using geographic-related features as pretext tasks is the key factor in appraising the price of real estate. Nam et al. (2023) designed self-generated tasks with pseudo-labels that have significant correlations with the downstream labels (e.g., predicting real estate prices through location and property size is similar to a new task that predicts rental rates by location and property size).
@@ 7 NS-TD Datasets and Benchmarks
In addition to application issues, fair dataset selection is also important for benchmarking different NS-TD algorithms. Several benchmarks proposed over the past few years have attempted to address one of the debatable concerns – Are deep learning approaches (including SSL) better than tree-based approaches? Due to the vast diversity of tabular data, algorithms that are state-of-the-art are significantly impacted by dataset selections and the corresponding hyperparameter tuning. To this end, recent NS-TD benchmarks for evaluating not only SSL4NS-TD but also machine learning as well as deep learning methods are presented in this section.
The commencement for standardized benchmarking was first introduced in DLBench Shwartz-Ziv and Armon (2022) and MLPCBench Kadra et al. (2021). DLBench compared deep learning algorithms against XGBoost Chen and Guestrin (2016) on 11 diverse tabular datasets with 7,000 to 1,000,000 samples, finding that XGBoost outperformed deep learning approaches Popov et al. (2020); Arik and Pfister (2021); Baosenguo (2021); Katzir et al. (2021) on most datasets, with no single deep learning algorithm consistently outperforming any other. MLPCBench evaluated a larger benchmark of 40 diverse tabular datasets with 400 to 400,000 samples, again illustrating that XGBoost outperformed deep learning approaches Erickson et al. (2020); Popov et al. (2020); Arik and Pfister (2021) on most datasets. However, given an improved hyperparameter optimization setup across regularization techniques, “cocktail” MLPs Kadra et al. (2021) can outperform both XGBoost and specialized deep learning approaches. DLBench and MLPCBench demonstrate the clear necessity for more robust approaches against existing works.
@@
Given the recent progress of SSL4NS-TD, TabularBench Grinsztajn et al. (2022) extended more comprehensive analyses including classification and regression tasks between tree-based models and SSL4NS-TD methods Gorishniy et al. (2021); Somepalli et al. (2022) across 45 medium-sized datasets with 3,000 to 10,000 samples. TabularBench revealed that tree-based models yield superior predictions more consistently with much less computational cost since SSL4NS-TD approaches suffer from oversmoothing tabular decision boundaries, noisy features, and misrepresenting non rotation-invariant data.
TabZilla McElfresh et al. (2023) conducted a large-scale study on 19 algorithms consisting of machine learning, deep learning, and SSL4NS-TD algorithms over 176 classification tabular datasets, and provided a benchmark of 36 datasets with 300 to 1,000,000 samples. TabZilla uncovered that CatBoost Prokhorenkova et al. (2018) achieved overall state-of-the-art performance, while TabPFN Hollmann et al. (2023) performed state-of-the-art performance on datasets with less than 1,250 samples. These evaluations demonstrate the diversity of tabular data and the need for tailored methods, and deep learning-based methods hinder their performance on more heterogeneous data. On the other hand, gradient-boosted decision trees perform better on irregular data, with skewed distribution or noisy features. Overall, XGBoost, TabPFN, and Catboost achieved the most consistent performance on the TabZilla benchmark.
Previous tabular learning benchmarks have demonstrated the achievements of standardizing datasets and training setups. More recently, TabPretNet Ye et al. (2023) released 1,000 labeled and 1,000 unlabeled datasets to standardize pretraining datasets Anonymous (2024a) for SSL4NS-TD. In addition to dataset selection, specifying the data scope tackled by new algorithms is important. For example, TabPFN Hollmann et al. (2023) serves as state-of-the-art performance only on small datasets, whereas Transformer-based approaches Gorishniy et al. (2021, 2022) specify their need for larger datasets. Lastly, XGBoost and Catboost still achieve consistent performance in most data regimes, indicating they are still competitive candidates to be compared against. These benchmarks offer unified evaluation collections that enable researchers to design new algorithms for diverse domains, and show that there remains potential to improve the current blind spots in the NS-TD problems.
@@ 8 Future Directions
Despite the appreciable achievements of SSL4NS-TD, there still remain unresolved challenges that need to be solved. Several future directions are listed for reference based on the existing literature.
@@ 8.1 A Recipe for SSL4NS-TD
Despite the existing explorations on various pretext tasks including predictive learning, contrastive learning, and hybrid learning, the SSL techniques of most existing works are mainly motivated by the success paradigms in NLP and CV. However, it is still unclear which SSL methods are more appropriate for the specific tabular scenarios and how to probe suitable hyper-parameters (e.g., the common masking ratio in the reconstruction pretext task and the proper batch size in contrastive learning), which are especially critical from the industrial perspective. While there is a potential research direction that lies in designing a pretext task that is correlated with the downstream applications Lee et al. (2021), it remains a challenging research problem.
@@ 8.2 Evolution of Foundation Tabular Models
Nowadays, foundation models such as ChatGPT have shown powerful dexterity in a variety of NLP applications, but foundation tabular models are unexplored mainly due to their heterogeneous, implicit, as well as order-invariant table characteristics. Although some recent works have undertaken the potential effectiveness of pre-training foundation tabular models from scratch (e.g., Yang et al. (2023)) and from PLMs (e.g., Anonymous (2024a)), there still exists much exploration space for a unified foundation model that is consistently superior to both deep learning and machine learning models. As tabular data play a vital role across diverse real-world fields, exploring foundation models for serving a diverse range of tabular applications remains an ongoing area of research.
8.3 More Effective and Efficient Methods
@@
Previous studies have demonstrated the effectiveness of developing deep learning models and adding SSL approaches across various tabular datasets. As reviewed in Section 7, tree-based models are still competitive or even better options because of their compelling performance but light computational cost in terms of many-shot scenarios. In addition, most prevailing literature centers on bolstering the performance of classification problems, while omitting more challenging yet critical regression problems, which do not have definitive boundaries of labels. With the advancements of incorporating larger scale tabular data and different modalities with distillation, we believe that the investigation on SSL4NS-TD for these perspectives can highlight future research on more robust yet deployable approaches.
@@ 9 Conclusion
SSL4NS-TD serves as a ubiquitous and vital connection between deep learning and applications with implicit relations. In this paper, we survey the existing SSL4NS-TD literature and provide an extensive review of advanced SSL4NS-TD training strategies, including predictive learning, contrastive learning, as well as hybrid learning. Three application issues are covered to showcase the promising potential of SSL4NS-TD. To facilitate reproducible research and compare the effectiveness of deep learning and tree-based models, we initiate the first step to summarize the representative NS-TD benchmarks and commonly used datasets for the research community. Furthermore, we highlight critical challenges and potential directions of SSL4NS-TD for future research. To the best of our knowledge, this is the first survey of SSL4NS-TD. We hope this survey can highlight the current research status of SSL4NS-TD in a unified view and shed light on future work on this promising paradigm.

### v3

@@ 6.2 Cross-Table Transferability
Directly learning representations from a table requires a trained model for each downstream dataset, and suffers from strict features between training and testing data, causing a severe burden in scaling and transferring to different problems. Therefore, how to learn representations across tables has been a critical demand for NS-TD systems in real-world scenarios, which benefits from reducing efforts on engineering features based on each dataset. Recent approaches achieving transferability involve SSL pre-training from PLMs to contextualize knowledge with coherent semantics (e.g., LIFT, TP-BERTa, GReaT) and from scratch with fine-grained feature encodings (e.g., TransTab, XTab, UniTabE). These methods have demonstrated that pre-training with SSL4NS-TD confers advantages for adaptation to incremental columns, low-resource scenarios, and missing value predictions [47, 53].
@@ 6.3 Domain Knowledge Integration
Tabular applications often have the need to incorporate expert knowledge to infer the results (e.g., clinical trials in the medical and real estate market in finance). [13] discovered that using geographic-related features as pretext tasks is the key factor in appraising the price of real estate. [30] designed self-generated tasks with pseudo-labels that have significant correlations with the downstream labels (e.g., predicting real estate prices through location and property size is similar to a new task that predicts rental rates by location and property size).
@@ 7.1 Existing NS-TD Datasets
In addition to application issues, fair dataset selection is also important for benchmarking different NS-TD algorithms. Due to the vast diversity of tabular data, state-of-the-art algorithms are significantly impacted by dataset selections and the corresponding hyperparameter tuning. To this end, recent NS-TD benchmarks demonstrate evaluating SSL4NS-TD against both machine learning and deep learning algorithms on datasets of varying properties, as summarized in Table 2.
Most NS-TD benchmarks are derived from the UCI Machine Learning Repository and the OpenML platform. Both sources feature a large collection of classification and regression NS-TD datasets. In particular, the OpenML platform released a benchmark suite, OpenML-CC18 [7], consisting of 72 curated classification NS-TD datasets. Despite pushes for standardization, most NS-TD and SSL4NS-TD algorithms focus on a subset of datasets from these 2 sources.
@@
The choice of datasets is important. Issues with dataset selection among NS-TD algorithms were first introduced in DLBench [34], which compared tree-based algorithms to recently proposed deep learning algorithms on 11 classification and regression datasets with 7,000 to 1,000,000 samples. These works found that XGBoost [10] outperformed more recent deep learning approaches [33, 1, 6, 22], when evaluated on their shared datasets. Hence, experimental results can easily be skewed by dataset selection. TabularBench [15] extended this analysis to SSL4NS-TD [14, 35] across 45 classification and regression task datasets with 3,000 to 10,000 samples. TabularBench revealed that tree-based models yield superior predictions more consistently with much less computational cost since both deep learning and SSL4NS-TD approaches suffer from oversmoothing tabular decision boundaries, noisy features, and misrepresenting non-rotation-invariant data. However, these conclusions mainly extend to medium-sized datasets.
TabZilla [29] conducted the most comprehensive large-scale study against 19 machine learning, deep learning [14], and SSL4NS-TD algorithms [35, 1, 50, 19] across 176 classification tabular datasets, and provided a benchmark of the 36 hardest datasets with 300 to 1,000,000 samples. Tabzilla uncovered none of the existing approaches perform well on all datasets. Particularly, transformer-based SSL4NS-TD algorithms [1] perform well on larger datasets while PFN SSL4NS-TD algorithms [19] perform well on smaller datasets. XGBoost [10] performs well on datasets with many irregular features. Among SSL4NS-TD algorithms, TabPFN [19] performs particularly well across the benchmark, definitively outperforming tree-based models with less than 1,250 samples. SAINT [35] comes in second across SSL4NS-TD algorithms on the TabZilla benchmark. Tabzilla paints a fairer picture of SSL4NS-TD, where specific algorithms excel under specific dataset properties.
Previous tabular learning benchmarks have demonstrated the achievements of standardizing datasets and training setups from OpenML and UCI datasets. While most existing SSL4NS-TD algorithms obtain datasets from a mix of the OpenML platform and UCI Machine Learning Repository, others target more specialized datasets. In particular, we found that several algorithms target vision datasets [38], synthetic datasets [23, 38, 8, 12], housing datasets [13], and medical datasets [38, 50, 40]. While standardized benchmarks are necessary to elucidate the data regimes each SSL4NS-TD algorithm is tailored for, application-based datasets show real-world use cases of the approach.
As the benefits of SSL4NS-TD become more apparent, effort has been placed on improving the pretraining dataset. While most SSLNS-TD algorithms perform standard train/dev/test dataset splitting, recent work curates larger and higher-quality datasets. OpenTabs [48] released around 2,000 unlabeled datasets. TP-BERTa [46, 48] released 202 labeled and 145 unlabeled datasets derived from OpenTabs. UniTabE [47] released 283,000 unlabeled datasets. The open-source effort to standardize pretraining datasets further accelerates research in SSL4NS-TD.
Algorithm Condorcet Rank
Wins $\uparrow$ Mean $\pm$ Std. $\downarrow$ Median $\downarrow$ Min $\downarrow$ Max $\downarrow$
VIME 3 3.800 $\pm$ 0.600 4.0 3.0 5.0
TabNet 26 2.700 $\pm$ 1.345 3.0 1.0 5.0
TabPFN 56 1.500 $\pm$ 0.671 1.0 1.0 3.0
TransTab 43 4.800 $\pm$ 0.400 5.0 4.0 5.0
SAINT 43 2.200 $\pm$ 0.600 2.0 1.0 3.0
Table 3: Performance across competitive approaches of each learning group. Rankings are aggregated across 36 datasets. The Condorcet winner is computed by counting the number of pair-wise wins each algorithm has across the benchmark.
7.2 Evaluations of SSL4NS-TD on the Common Benchmark
Several popular SSL tabular learning algorithms from each learning category are benchmarked on the most recent TabZilla benchmark [29] consisting of 36 datasets. Specifically, we experimented each algorithm for 5 optional hyperparameter optimization steps with 10-fold cross-validation under the same experimental settings as TabZilla. We set the time limit across all validation splits and hyperparameter tunings to 2 hours to reflect limited computational resources. As not all algorithms complete in 2 hours, we report the rankings of the algorithms across a shared set of datasets, following TabZilla. Because this evaluation metric drops datasets with incomplete relative rankings [45], we also evaluate by Condorcet voting [39], which ranks models based on pair-wise ranking scores, extracting the maximum information from datasets with incomplete rankings.
As summarized in Table 3, TabPFN achieves the best average performance among SSL4NS-TD algorithms, followed by SAINT, TabNet, VIME, and TransTab. However, the worst ranks of TabPFN and SAINT are the third rank, while TabNet is able to achieve the best performance in at least one dataset. Hence, the optimal SSL4NS-TD algorithm remains dataset-dependent, which also indicates that there is no state-of-the-art SSL4NS-TD approach that is superior across all datasets. Nonetheless, TabPFN receives the overall best ranks and wins, suggesting that TabPFN serves as an appropriate initial choice for consideration. The comparison between SAINT and TabPFN illustrates that integrating predictive learning and contrastive learning may need to be further investigated.
@@ 8 Future Directions
Witnessing the appreciable achievements of SSL4NS-TD, we have foreseen unresolved challenges based on the literature and the demands of real-world scenarios. In this section, we identify and discuss key research opportunities that merit future exploration.
@@ 8.1 A Recipe for SSL4NS-TD
Despite the existing explorations on various pretext tasks including predictive learning, contrastive learning, and hybrid learning, the SSL techniques of most existing works are mainly motivated by the success paradigms in NLP and CV. However, it is still unclear which SSL methods are more appropriate for the specific tabular scenarios and how to probe suitable hyper-parameters (e.g., the common masking ratio in the reconstruction pretext task and the proper batch size in contrastive learning), which are especially critical from the industrial perspective. While there is a potential research direction that lies in designing a pretext task that is correlated with the downstream applications [24], more research efforts are needed to address the above problems.
@@ 8.2 Evolution of Foundation Tabular Models
Representation learning is a fundamental yet crucial task in tabular domains, particularly in the realm of NS-TD datasets. Recently, foundation models such as ChatGPT have shown powerful dexterity in a variety of NLP applications, but foundation tabular models are unexplored mainly due to their heterogeneous, implicit, as well as order-invariant table characteristics. Although some recent works have undertaken the potential effectiveness of pre-training foundation tabular models from scratch (e.g., [47]) and from PLMs (e.g., [46]), there still exists much exploration space for a unified foundation model that is consistently superior to both deep learning and machine learning models. As tabular data play a vital role across diverse real-world fields, exploring foundation models for serving a diverse range of tabular applications remains an ongoing area of research.
8.3 Continual Learning with Tabular Models
While the transferability of LLMs has been showcased in improving downstream performance, this approach remains significantly underexplored in the context of training foundation models for non-sequential tabular data. The challenges arise due to the immature developments of foundation tabular models as well as the diverse and heterogeneous characteristics across different tabular datasets. A potential solution is to unify the format (e.g., JSON) of all tabular datasets to be consumed by the foundation tabular models or existing LLMs. Nonetheless, the development of continual learning strategies for tabular models has emerged as a promising avenue for future research, with the potential to reduce resource consumption and improve the adaptation of these models to current information.
8.4 Privacy-Aware Tabular Models
While learning from a large volume of tabular data benefits learning representations, it is challenging to centrally train a foundation tabular model due to the sensitive and private proprietary nature of tabular applications (e.g., financial and healthcare stakeholders). One of the potential approaches is to leverage federated learning, where multiple devices or servers collaboratively train a model without sharing their local data [37]. Instead of centralizing the data, each participating stakeholder trains the model on its own local dataset and shares only the model updates (such as gradients or weights) with a central server. The server then aggregates these updates to improve the global model. Nevertheless, the direct usage of federated learning in tabular applications yields challenges, primarily stemming from data quality and distributions of tabular data. Addressing these substantial issues opens up a challenge, thereby establishing a promising research direction for the application of federated learning in tabular domains.
8.5 Advancements in Multi-Modal Multi-Task Environments
@@
While previous studies have demonstrated the effectiveness of developing deep learning models and adding SSL approaches across various tabular datasets, as reviewed in Section 7, tree-based models are still competitive or even better options because of their compelling performance but light computational cost in terms of many-shot scenarios. In addition, regression problems are critical in the NS-TD applications, e.g., health metrics in the medical domain and house price prediction in the financial domain. However, most prevailing literature centers on bolstering the performance of classification problems, while omitting more challenging yet critical regression problems, which do not have definitive boundaries of labels. Supporting multi-task abilities may augment tabular models to share common knowledge for different tasks while providing memory-efficient resources to host them.
In addition, most existing NS-TD methods have focused on tabular-based information. However, the multi-modal integration has demonstrated effectiveness in the sequential tabular domains, e.g., training with images of the items and the corresponding metadata for recommendations [41]. Furthermore, the conversion from tabular format to text format (e.g., 0 $\rightarrow$ seen; 1 $\rightarrow$ unseen) enables the use of leveraging generic-knowledge LLMs to learn rich context. With the advancements of incorporating larger scale tabular data and different modalities, we believe that the investigation on SSL4NS-TD for these perspectives can highlight future research on more robust yet deployable approaches.
@@ 9 Conclusion
SSL4NS-TD serves as a ubiquitous and vital connection between deep learning and applications with implicit relations. In this paper, we survey the existing SSL4NS-TD literature and provide an extensive review of advanced SSL4NS-TD training strategies, including predictive learning, contrastive learning, as well as hybrid learning. Three application issues are covered to showcase the promising potential of SSL4NS-TD. To facilitate reproducible research and compare the effectiveness of deep learning and tree-based models, we initiate the first step to summarize the representative NS-TD benchmarks and commonly used datasets for the research community, with the results of multiple recent SSL4NS-TD approaches on the large-scale benchmark. Furthermore, we highlight critical challenges and potential directions of SSL4NS-TD for future research. To the best of our knowledge, this is the first survey of SSL4NS-TD. We hope this survey can highlight the current research status of SSL4NS-TD from a unified view and shed light on future work on this promising paradigm.